OpenStudy (nathaniel):

Please help me! Consider the quadratic function f(x)=x^2 - 8x +195 Write in Vertex form. Identify the vertex. Equation line of symmetry. Solve the Equation. What are the x Intercepts?

OpenStudy (rida34):

Do you mean the axis of symmetry?

OpenStudy (rida34):

either way, I'll try to explain: 1. To go from standard form to vertex, you need to complete the square - it's hard to explain but I suggest watching a YouTube video on it since once you get it it's fairly simple. Basically you take your second term, divide it by 2, square it and then add and subtract in order to balance your equation ( sorry this is not the best explanation) = x^2 - 8x + 195 = (x^2 - 8x +16 - 16) + 195 - you'll recognize that x^2 - 8x + 16 is a perfect square = (x-4)^2 - 16 + 195 - the number that's not part of the perfect square, you take it out of the bracket = (x-4)^2 - 179 - your vertex is (4, -179) 2. The axis of symmetry is the x value of the coordinate, in this case it would be 4. 3. I'm not sure what you mean by solving it, maybe try using the quadratic formula to find the x values? 4. To find the x intercepts, factor it